the half life for a first order reaction is 25 minutes. calculate the rate constant and apply it ti find the time required for 85% complition of the reaction ?

Grade:12th pass

1 Answers

Sunil Kumar FP
9 years ago
we have the relation between the half lide and the rate constant as
half life t1/2=.693/k
k is the rate constant
or k=.693/t1/2
=.693/25 min^-1
=2.772*10^-2 min^-1

time required for 25% completition
k=2.303/t log (initial conc/final conc)
2.772*10^-2=2.303/t log (100/15)
t=68.3 min